OpenStudy (anonymous):

For the polynomial 3x^3+4x^2-7x+2=0, are there any other rational roots other than 2/3? When I try to factor 3x^2+6x-3, I get nowhere.

OpenStudy (anonymous):

\[ \frac23\ |\ 3\ 4\ -7\ \ \ \ \ 2\\ \ \ \ \ \ \ \ \ \ 2\ \ \ \ \ 4\ -2\\ \ \ \ \ \ \ 3\ 6\ -3\ \ \ \ \ 0 \\ (x-\frac23)(3x^2+6x-3)=(3x-2)(x^2+2x-1)=0\\ x^2+2x-1=0\\ x=\frac{-2\pm2\sqrt2}2=-1\pm\sqrt2 \]

OpenStudy (anonymous):

So, no, there are no more rational roots... the other 2 are irrational.

OpenStudy (anonymous):

Okay. Thanks